# Deformation quantization and perverse sheaves

### v1

@@ 4.3. Tensor product of DQ modules
Consider two objects $\mathcal{L},\mathcal{M}\in\mathrm{RH}_{S}$ which are regular holonomic $\rho_{*}\mathsf{E}_{Y}$-modules. In this section we will study the complex of $\mathbf{C}(\!(\hbar)\!)$-linear sheaves $\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$ on $X$. Our first goal is to endow it with extra structure.
Choose a contactification $\rho\colon Y\rightarrow S$ and assume that $\mathsf{E}_{Y}$ and $\mathsf{W}_{S}$ are represented by sheaves of algebras, there is an element $\hbar\in\mathsf{E}_{Y}(-1)$ with principal symbol $\tau^{-1}$ satisfying $\hbar^{*}=-\hbar$ and $\mathsf{W}_{S}$ is the subalgebra of $\rho_{*}\mathsf{E}_{Y}$ of elements commuting with $\hbar$.
@@
$\mathsf{B}_{Y}=(\rho_{*}\mathsf{E}_{Y})\otimes_{\mathsf{W}_{S}}(\rho_{*}\mathsf{E}_{Y}),$
where the right action of $\mathsf{W}_{S}$ on the first copy of $\rho_{*}\mathsf{E}_{Y}$ is via $P\in\mathsf{W}_{S},Q\in\rho_{*}\mathsf{E}_{Y}\mapsto P^{*}Q$. It carries a natural bimodule structure as follows: $\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y}$ acts in an obvious way on the right and $\hbar$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $Q_{1}\otimes\hbar Q_{2}=-\hbar Q_{1}\otimes Q_{2}$.
@@
$\mathsf{B}_{Y}$ has a canonical structure of a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y})$-bimodule, where $D\hbar=\hbar(D+1)$.
@@
Let $\partial_{t}=\hbar^{-1}\in\mathsf{E}_{Y}(1)$. Locally we may choose an element $t\in\mathsf{E}_{Y}(0)$ satisfying
@@
This allows us to enhance $\mathsf{B}_{Y}$ to a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y})$-bimodule, where $D=\partial_{t}t$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $-D^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes DQ_{2}$.
Let us show that this enhancement is independent of the choice of $t$. If $t^{\prime}$ is another element of $\mathsf{E}_{Y}(0)$ satisfying $[\partial_{t},t^{\prime}]=1$, then $[\partial_{t},t-t^{\prime}]=0$, i.e. $[\hbar,t-t^{\prime}]=0$. Thus, there is an element $R\in\mathsf{W}_{S}(0)$ such that
@@
But $\partial_{t}R\in\mathsf{W}_{S}$ and hence the last two terms cancel. ∎
@@
For any $\mathcal{L},\mathcal{M}\in\mathrm{RH}_{S}$ the sheaf $\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$ has a canonical structure of a differential perverse sheaf.
@@
Consider the duality functor $\mathbb{D}(\mathcal{L})=\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathcal{L},\mathsf{W}_{S})[\dim S/2]$. By [KS12, Lemma 7.2.2] $\mathbb{D}(\mathcal{L})$ is a holonomic $\mathsf{W}_{S}$-module concentrated in degree 0 and we have
$\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]\cong\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}.$
It is then shown in [KS12, Theorem 7.2.3] that $\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]$ is a perverse sheaf of $\mathbf{C}(\!(\hbar)\!)$-vector spaces.
$\rho_{*}\mathsf{E}_{Y}$ is flat as a $\mathsf{W}_{S}$-module by Lemma 4.4. Therefore,
$\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}\cong\mathsf{B}_{Y}\otimes^{\mathbb{L}}_{\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y}}(\mathcal{L}\otimes\mathcal{M})$
which has a differential structure by 4.11. ∎
@@
Consider two regular holonomic modules $\mathcal{L},\mathcal{M}\in\mathrm{RH}_{S}$ and assume $\mathcal{L}$ is simple along a Lagrangian submanifold $L\subset S$. Then
$\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$
@@
To define the differential structure we may choose $t\in\mathsf{E}_{Y}(0)$ whose principal symbol $\sigma_{0}(t)\in\mathcal{O}_{Y}$ vanishes along $\Lambda$. With this choice $D=\partial_{t}t=\hbar\partial_{\hbar}\in\mathsf{E}_{\Lambda/Y}$. By assumption $\mathcal{L}$ is an $\mathsf{E}_{Y}$-module simple along $\Lambda$. Thus, there is an $\mathsf{E}_{\Lambda/Y}$-lattice $\mathcal{L}_{0}\subset\mathcal{L}$ generating $\mathcal{L}$ over $\mathsf{E}_{Y}$ and $D$ acts on
¯$\overline{\mathcal{L}}=\mathcal{L}_{0}/\mathsf{E}_{\Lambda/Y}(-1)\mathcal{L}_{0}$
@@
Let $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ be a $\mathsf{E}_{\Lambda/Y}$-lattice such that no two eigenvalues of $D$ on
¯$\overline{\mathcal{M}}=\mathcal{M}_{0}/\mathsf{E}_{\Lambda/Y}(-1)\mathcal{M}_{0}$
@@
By [KS12, Lemma 7.1.3]
$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\mathsf{W}_{L/S}}\mathcal{M}_{0}$
is a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$. Since $D\in\mathsf{E}_{\Lambda/Y}$, this lattice is preserved by $D$. Moreover, no two eigenvalues of $D$ on the $\mathbf{C}$-linear perverse sheaf
$(\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\mathsf{W}_{L/S}}\mathcal{M}_{0})/\hbar\cong\overline{\mathcal{L}}_{0}\otimes_{\mathsf{W}_{L/S}/\hbar}\overline{\mathcal{M}}_{0}$
@@
We will now give a description of the above differential perverse sheaf using microlocalization. Consider the setting of 4.13 and suppose $L\subset S$ is lifted to a Legendrian $\Lambda\subset Y$ in a contactification $\rho\colon Y\rightarrow S$. In this case we may identify $\tilde{\Lambda}\cong L\times\mathbf{C}^{\times}$. Then $\mu^{dR}_{\Lambda}\mathcal{L},\mu^{dR}_{\Lambda}\mathcal{M}$ are twisted $D$-modules on $\tilde{\Lambda}$, monodromic with respect to the $\mathbf{C}^{\times}$-action. Therefore,
@@
is a perverse sheaf on $\tilde{\Lambda}=L\times\mathbf{C}^{\times}$, lisse in the $\mathbf{C}^{\times}$-direction. Its restriction to $L\times\{1\}$ defines a perverse sheaf
@@
Consider the setting of 4.13. There is an isomorphism of differential perverse sheaves
$\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}\cong\mathrm{RH}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{M})|_{L}[-1]\right)$
@@
As in the proof of 4.13 we choose $\mathsf{E}_{\Lambda/Y}$-submodules $\mathcal{L}_{0}\subset\mathcal{L}$ and $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ generating these modules over $\mathsf{E}_{Y}|_{\Lambda}$ as well as an element $D=\partial_{t}t\in\mathsf{E}_{\Lambda/Y}$.
@@
$\operatorname{gr}^{0}_{\Lambda}\mathsf{E}_{\Lambda/Y}\cong\mathcal{D}^{\sqrt{v}}_{L}[\theta],$
@@
$\mathsf{E}_{\Lambda/Y}\longrightarrow\operatorname{gr}^{0}_{\Lambda}\mathsf{E}_{\Lambda/Y}.$
@@
As $\hbar\mathsf{E}_{\Lambda/Y}(0)=\mathsf{E}_{\Lambda/Y}(-1)$, these are exactly the associated graded sheaves with respect to the microlocal $V$-filtration
@@
$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\mathsf{W}_{L/S}}\mathcal{M}_{0}$
defines a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes_{\mathsf{W}_{S}}\mathcal{M}$ preserved by $D$. Therefore, by 2.11 the claim will follow once we establish an isomorphism
@@
given by $Q_{1}\otimes Q_{2}\mapsto Q_{2}^{\prime}\otimes Q_{1}^{\prime}$ of $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y})$-bimodules.
For an object $\mathcal{M}\in\mathrm{RH}_{\mathrm{T}^{*}X}$ denote by $\mathcal{M}^{\prime}$ the same module with the ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$-action twisted by the involution $(-)^{\prime}$. Then the above involution of $\mathsf{B}_{Y}$ induces an isomorphism
@@ 5.1. Vanishing cycles
Let $X$ be a complex manifold equipped with a holomorphic function $f\colon X\rightarrow\mathbf{C}$. Recall the perverse sheaf $\phi_{f}\in\mathrm{Perv}(f^{-1}(0);\mathbf{C})$ of vanishing cycles of the constant perverse sheaf $\mathbf{C}_{X}[\dim X]\in\mathrm{Perv}(X)$ [Dim04], which carries a monodromy automorphism $T$. In this section we give a microlocal description of this sheaf.
@@
The standard projection $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ provides a contactification, so we may regard $\mathcal{C}^{\lambda,\sqrt{X}}_{f}$ as an object of $\mathrm{RH}_{\mathrm{T}^{*}X}$. This is a simple module along the Lagrangian $\Gamma_{df}\subset\mathrm{T}^{*}X$ of order $\lambda$.
@@
The following statement is due to Kashiwara and Malgrange, see [Mal83, Théorème 6.2.6] and [Kas83, Theorem 2].
@@
Using this microlocal perspective we can prove a relationship between vanishing cycles and DQ modules. Recall from 4.13 that
@@
$M_{f}\colon\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\widehat{\mathrm{RH}}^{-1}(\phi_{f}\otimes\mathbf{C}_{\mu+\lambda}),$
@@
Consider the zero section $L=X\subset\mathrm{T}^{*}X$ and its contactification $\Lambda=j^{1}0\subset\mathrm{J}^{1}X$. $\mathcal{C}_{0}$ as a $\widehat{\mathcal{W}}_{X}$-module is simple along $L$. Using 4.14 we get an isomorphism
$\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\widehat{\mathrm{RH}}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\mathcal{D}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{C}^{\mu}_{f})|_{L}[-1]\right)$

### v2

@@ 4.3. Tensor product of DQ modules
Consider two regular holonomic $\widehat{\mathsf{E}}_{Y}$-modules $\mathcal{L},\mathcal{M}$. In this section we will study the complex of $\mathbf{C}(\!(\hbar)\!)$-linear sheaves $\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$ on $X$. Our first goal is to endow it with extra structure.
Choose a contactification $\rho\colon Y\rightarrow S$ and assume that $\widehat{\mathsf{E}}_{Y}$ and $\widehat{\mathsf{W}}_{S}$ are represented by sheaves of algebras, there is an element $\hbar\in\widehat{\mathsf{E}}_{Y}(-1)$ with principal symbol $\tau^{-1}$ satisfying $\hbar^{*}=-\hbar$ and $\widehat{\mathsf{W}}_{S}$ is the subalgebra of $\rho_{*}\widehat{\mathsf{E}}_{Y}$ of elements commuting with $\hbar$.
@@
$\mathsf{B}_{Y}=(\rho_{*}\widehat{\mathsf{E}}_{Y})\otimes_{\widehat{\mathsf{W}}_{S}}(\rho_{*}\widehat{\mathsf{E}}_{Y}),$
where the right action of $\widehat{\mathsf{W}}_{S}$ on the first copy of $\rho_{*}\widehat{\mathsf{E}}_{Y}$ is via $P\in\widehat{\mathsf{W}}_{S},Q\in\rho_{*}\widehat{\mathsf{E}}_{Y}\mapsto P^{*}Q$. It carries a natural bimodule structure as follows: $\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y}$ acts in an obvious way on the right and $\hbar$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $Q_{1}\otimes\hbar Q_{2}=-\hbar Q_{1}\otimes Q_{2}$.
@@
$\mathsf{B}_{Y}$ has a canonical structure of a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y})$-bimodule, where $D\hbar=\hbar(D+1)$.
@@
Let $\partial_{t}=\hbar^{-1}\in\widehat{\mathsf{E}}_{Y}(1)$. Locally we may choose an element $t\in\widehat{\mathsf{E}}_{Y}(0)$ satisfying
@@
This allows us to enhance $\mathsf{B}_{Y}$ to a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y})$-bimodule, where $D=\partial_{t}t$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $-D^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes DQ_{2}$.
Let us show that this enhancement is independent of the choice of $t$. If $t^{\prime}$ is another element of $\widehat{\mathsf{E}}_{Y}(0)$ satisfying $[\partial_{t},t^{\prime}]=1$, then $[\partial_{t},t-t^{\prime}]=0$, i.e. $[\hbar,t-t^{\prime}]=0$. Thus, there is an element $R\in\widehat{\mathsf{W}}_{S}(0)$ such that
@@
But $\partial_{t}R\in\widehat{\mathsf{W}}_{S}$ and hence the last two terms cancel. ∎
@@
For any regular holonomic $\rho_{*}\widehat{\mathsf{E}}_{Y}$-modules $\mathcal{L},\mathcal{M}$ the sheaf $\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$ has a canonical structure of a differential perverse sheaf.
@@
Consider the duality functor $\mathbb{D}(\mathcal{L})=\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathcal{L},\widehat{\mathsf{W}}_{S})[\dim S/2]$. By [KashiwaraSchapira, Lemma 7.2.2] $\mathbb{D}(\mathcal{L})$ is a holonomic $\widehat{\mathsf{W}}_{S}$-module concentrated in degree 0 and we have
$\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]\cong\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}.$
It is then shown in [KashiwaraSchapira, Theorem 7.2.3] that $\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]$ is a perverse sheaf of $\mathbf{C}(\!(\hbar)\!)$-vector spaces.
$\rho_{*}\widehat{\mathsf{E}}_{Y}$ is flat as a $\widehat{\mathsf{W}}_{S}$-module by Lemma 4.4. Therefore,
$\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}\cong\mathsf{B}_{Y}\otimes^{\mathbb{L}}_{\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y}}(\mathcal{L}\otimes\mathcal{M})$
which has a differential structure by 4.10. ∎
@@
Consider two regular holonomic $\rho_{*}\widehat{\mathsf{E}}_{Y}$-modules and assume $\mathcal{L}$ is simple along a Lagrangian submanifold $L\subset S$. Then
$\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$
@@
To define the differential structure we may choose $t\in\widehat{\mathsf{E}}_{Y}(0)$ whose principal symbol $\sigma_{0}(t)\in\mathcal{O}_{Y}$ vanishes along $\Lambda$. With this choice $D=\partial_{t}t=\hbar\partial_{\hbar}\in\widehat{\mathsf{E}}_{\Lambda/Y}$. By assumption $\mathcal{L}$ is an $\widehat{\mathsf{E}}_{Y}$-module simple along $\Lambda$. Thus, there is an $\widehat{\mathsf{E}}_{\Lambda/Y}$-lattice $\mathcal{L}_{0}\subset\mathcal{L}$ generating $\mathcal{L}$ over $\widehat{\mathsf{E}}_{Y}$ and $D$ acts on
¯$\overline{\mathcal{L}}=\mathcal{L}_{0}/\widehat{\mathsf{E}}_{\Lambda/Y}(-1)\mathcal{L}_{0}$
@@
Let $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ be a $\widehat{\mathsf{E}}_{\Lambda/Y}$-lattice such that no two eigenvalues of $D$ on
¯$\overline{\mathcal{M}}=\mathcal{M}_{0}/\widehat{\mathsf{E}}_{\Lambda/Y}(-1)\mathcal{M}_{0}$
@@
By [KashiwaraSchapira, Lemma 7.1.3]
$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{L/S}}\mathcal{M}_{0}$
is a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$. Since $D\in\widehat{\mathsf{E}}_{\Lambda/Y}$, this lattice is preserved by $D$. Moreover, no two eigenvalues of $D$ on the $\mathbf{C}$-linear perverse sheaf
$(\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{L/S}}\mathcal{M}_{0})/\hbar\cong\overline{\mathcal{L}}_{0}\otimes_{\widehat{\mathsf{W}}_{L/S}/\hbar}\overline{\mathcal{M}}_{0}$
@@
We will now give a description of the above differential perverse sheaf using microlocalization. Consider the setting of 4.12 and suppose $L\subset S$ is lifted to a Legendrian $\Lambda\subset Y$ in a contactification $\rho\colon Y\rightarrow S$. In this case we may identify $\tilde{\Lambda}\cong L\times\mathbf{C}^{\times}$. Then $\mu^{dR}_{\Lambda}\mathcal{L},\mu^{dR}_{\Lambda}\mathcal{M}$ are twisted $D$-modules on $\tilde{\Lambda}$, monodromic with respect to the $\mathbf{C}^{\times}$-action. Therefore,
@@
is a perverse sheaf on $\tilde{\Lambda}=L\times\mathbf{C}^{\times}$, which is locally constant along the $\mathbf{C}^{\times}$ orbits. Its restriction to $L\times\{1\}$ defines a perverse sheaf
@@
Consider the setting of 4.12. There is an isomorphism of differential perverse sheaves
$\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}\cong\mathrm{RH}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{M})|_{L}[-1]\right)$
@@
As in the proof of 4.12 we choose $\widehat{\mathsf{E}}_{\Lambda/Y}$-submodules $\mathcal{L}_{0}\subset\mathcal{L}$ and $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ generating these modules over $\widehat{\mathsf{E}}_{Y}|_{\Lambda}$ as well as an element $D=\partial_{t}t\in\widehat{\mathsf{E}}_{\Lambda/Y}$.
@@
$\operatorname{gr}^{0}_{\Lambda}\widehat{\mathsf{E}}_{\Lambda/Y}\cong\mathcal{D}^{\sqrt{v}}_{L}[\theta],$
@@
$\widehat{\mathsf{E}}_{\Lambda/Y}\longrightarrow\operatorname{gr}^{0}_{\Lambda}\widehat{\mathsf{E}}_{\Lambda/Y}.$
@@
As $\hbar\widehat{\mathsf{E}}_{\Lambda/Y}(0)=\widehat{\mathsf{E}}_{\Lambda/Y}(-1)$, these are exactly the associated graded sheaves with respect to the microlocal $V$-filtration
@@
$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{L/S}}\mathcal{M}_{0}$
defines a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$ preserved by $D$. Therefore, by 2.12 the claim will follow once we establish an isomorphism
@@
given by $Q_{1}\otimes Q_{2}\mapsto Q_{2}^{\prime}\otimes Q_{1}^{\prime}$ of $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y})$-bimodules.
For a regular holonomic ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$-module $\mathcal{M}$ denote by $\mathcal{M}^{\prime}$ the same module with the ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$-action twisted by the involution $(-)^{\prime}$. Then the above involution of $\mathsf{B}_{Y}$ induces an isomorphism
@@ 5.1. Vanishing cycles
Let $X$ be a complex manifold equipped with a holomorphic function $f\colon X\rightarrow\mathbf{C}$. Recall the perverse sheaf $\phi_{f}\in\mathrm{Perv}(f^{-1}(0);\mathbf{C})$ of vanishing cycles of the constant perverse sheaf $\mathbf{C}_{X}[\dim X]\in\mathrm{Perv}(X)$ [Dimca], which carries a monodromy automorphism $T$. In this section we give a microlocal description of this sheaf.
@@
The following statement is due to Kashiwara and Malgrange, see [Malgrange, Théorème 6.2.6] and [KashiwaraVanishing, Theorem 2].
@@
Using this microlocal perspective we can prove a relationship between vanishing cycles and DQ modules. Recall from 4.12 that
@@
$M_{f}\colon\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\mathrm{RH}^{-1}(\phi_{f}\otimes\mathbf{C}_{\mu+\lambda}),$
@@
Consider the zero section $L=X\subset\mathrm{T}^{*}X$ and its contactification $\Lambda=j^{1}0\subset\mathrm{J}^{1}X$. $\mathcal{C}_{0}$ as a $\widehat{\mathcal{W}}_{X}$-module is simple along $L$. Using 4.13 we get an isomorphism
$\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\mathrm{RH}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\mathcal{D}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{C}^{\mu}_{f})|_{L}[-1]\right)$